After retiring, you want to be able to take $1000 every month for a total of 20 years from your retirement account. The account earns 6% interest. How much will you need in your account when you retire?


1
Expert's answer
2021-10-06T00:29:21-0400

This question requires us to determine the principal(starting amount, principal). To find its value, the payout annuity formula given below is used.

Payout annuity formula,

"P_0=m(1-(1+r\/n)^{-Z*(n)})\/(r\/n)" where,

"P_0" is the balance in the account at the beginning(starting amount , principal)

"m" is the regular withdrawal

"r" is the annual interest (decimal form)

"n" is the number of compounding periods in one year

"Z" is the number of years we plan to take withdrawals.

From the information above, we are given,

"m=1000"

"r=6\%=0.06"

"n=12"

"Z=20"

Substituting this in the formula,

"P_0=1000*(1-(1+0.06\/12)^ {(-20*(12)})\/(0.06\/12)"

"P_0=1000*(1-(1.005)^{-240})\/0.005"

"P_0=1000*(1-0.3020961)\/0.005"

"P_0=1000*0.6979039\/0.005=697.9039\/0.005=139580.80"

Therefore, I will need $139,580.80 in my account when I retire.